OpenStudy (anonymous):

Evaluate the integral using the properties of even and odd functions as an aid.

OpenStudy (tkhunny):

Well, do it. Is it even or odd? Is the range of integration symmetrical about the origin?

OpenStudy (anonymous):

I'm pretty sure it's odd.

OpenStudy (tkhunny):

Please write the definition of an Odd Function. Pretty sure isn't good enough,

OpenStudy (tkhunny):

Even: f(x) = f(-x) Is that true for this function? Odd: f(x) = -f(-x) Is that true for this function?

OpenStudy (anonymous):

\[\int\limits_{-2}^{2} x ^{2}(x ^{2}+1) dx\]

OpenStudy (anonymous):

Help please

OpenStudy (tkhunny):

You haven't answer the question. Is it Odd or Even? Check out those definitions.

OpenStudy (anonymous):

Oh its even

OpenStudy (tkhunny):

How do you know?

OpenStudy (anonymous):

An even function is a function such that f(-x)=f(x). Graphically an even function is symmetrical about the y-axis while an odd function is symmetrical when rotated about the origin 180 degrees.

OpenStudy (anonymous):

Am I right?

OpenStudy (tkhunny):

Let's see: \(f(-x) = (-x)^{2}((-x)^{2}+ 1) = x^{2}(x^{2}+1)= f(x)\) -- Perfect. Okay, how do we utilize this to simplify the integral?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

do you use\[u = x ^{2}+1 \] which is du is 2x

OpenStudy (anonymous):

while dx is du/2x

OpenStudy (tkhunny):

How does that utilize the fact that it is an Even Function? First, there is no need for a substitution. This is a simple integral. The lesson, here, is about symmetry.

OpenStudy (anonymous):

Im confused?

OpenStudy (tkhunny):

The point you are to see is that we have this situation: 1) Even Function 2) Symmetric About the y-Axis 3) Half the integral is redundant. \(\int\limits_{-2}^{2}x^{2}(x^{2}+ 1)\;dx = 2\cdot\int\limits_{0}^{2}x^{2}(x^{2}+ 1)\;dx\)

OpenStudy (anonymous):

oh yeah.. Now I got it

OpenStudy (anonymous):

after this do you add 1

OpenStudy (anonymous):

Sorry, mom called me

OpenStudy (tkhunny):

Add 1 for what? Just evaluate the integral.

OpenStudy (anonymous):

okay, I'll let me work on it as soon as possible. I'm eating dinner

OpenStudy (tkhunny):

I had Frog Legs and Pork Chops. It was pretty good.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

That's cool

OpenStudy (anonymous):

I don't eat meat

OpenStudy (tkhunny):

Also cool.

OpenStudy (anonymous):

I heard people eat all sorts of thing. None of them make me scared

OpenStudy (tkhunny):

Well, let's eat some calculus, shall we?

OpenStudy (anonymous):

lol Im almost done ... let me drink water

OpenStudy (tkhunny):

My frogs would like that.

OpenStudy (anonymous):

I find it little hard writing the equation down

OpenStudy (anonymous):

\[2 \int\limits_{0}^{2} x^{2}(x ^{2}+1)dx \]

OpenStudy (anonymous):

\[x ^{2} (x ^{2}+1) \int\limits_{0}^{5}\]

OpenStudy (anonymous):

Is this something like this

OpenStudy (tkhunny):

Remember, this is a trivial integral. If you are thinking of something fancy, you don't see it.

OpenStudy (anonymous):

wouldn't you subtitute (2^2 +1) = 5 and 0^2+ 1 which the bottom is 1

OpenStudy (anonymous):

instead of 0. If I'm wrong help?

OpenStudy (anonymous):

Am I right?

OpenStudy (tkhunny):

Umm... You should first find the anti-derivative. \(\int x^{2}(x^{2} + 1)\;dx = \int x^{4} + x^{2}\;dx = ??\)

OpenStudy (anonymous):

\[\frac{ x }{ 5}^{5}+ x ^{2}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I mean x^ 3

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }x ^{5}+ x ^{3}\]

OpenStudy (tkhunny):

Why didn't we get \(\dfrac{1}{5}x^{5} + \dfrac{1}{3}x^{3} + C\)

OpenStudy (anonymous):

oh yeah plus C

OpenStudy (anonymous):

I think you made a mistake its x cube

OpenStudy (anonymous):

never mind sorry

OpenStudy (anonymous):

What do you do after this? do you subtitute the 2 and 0 to get the answer for this question

OpenStudy (anonymous):

meanwhile you subtract both of them right

OpenStudy (tkhunny):

First, you get the integral right. If you think there is no 1/3 in there, you should review your notes. Why did you get a 1/5 on the first piece, but nothing on the second piece? That makes no sense.

OpenStudy (anonymous):

Well sorry.. wasn't paying attention.

OpenStudy (anonymous):

I don't know what is wrong with me

OpenStudy (anonymous):

I hope you understand me.. I make lot of mistake..

OpenStudy (anonymous):

I need to get this done because I have headache.. hope you understand right

OpenStudy (anonymous):

I just had seizure recently..

OpenStudy (tkhunny):

It is okay to make mistakes so long as you learn from them. \(2\cdot\int\limits_{0}^{2} x^{4} + x^{2}\;dx = \dfrac{1}{5}x^{5} + \dfrac{1}{3}x^{3}|_{0}^{2} = 2*(\left((1/5)32+(1/3)8\right)-\left(0\right)) = 2(35/5 + 8/3)\)

OpenStudy (anonymous):

did you got 32 by 3^5

OpenStudy (anonymous):

sorry 2^5

OpenStudy (anonymous):

I caught up

OpenStudy (anonymous):

is it 9\[9 \frac{ 2 }{ 3 }\]

OpenStudy (tkhunny):

?? Sorry, typo up there. 32, not 35 2(32/5 + 8/3) = 16(4/5 + 1/3) = 16(12/15 + 5/15) = 16 (17/15) It's just arithmetic.

OpenStudy (anonymous):

Am I right to the final answer?

OpenStudy (tkhunny):

You should be able to compare my correct response with yours. You tell me if yours is correct.

OpenStudy (anonymous):

it's wrong... but I should be multiplying by 2

OpenStudy (anonymous):

in the end

OpenStudy (anonymous):

Thanks for help!

OpenStudy (tkhunny):

Did you follow mine at all? The only error was that goofey 35 and the cut off of the final /3. Down below, all the arithmetic is shown. Keep trying until you get 16 (17/15) = 272/15

OpenStudy (anonymous):

alright I got it I did it.

OpenStudy (anonymous):

Thanks for help!